OpenStudy (anonymous):

How many solutions are there for the equation? x2 + 16 = 8x A. 2 B. 1 C. 0 I put B correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

:P Are you just guessing Satellite? :P Ur answering rly fast.. lol

OpenStudy (anonymous):

\[x^2-8x+16=(x-4)^2\]

OpenStudy (anonymous):

so only one answer: x=4

OpenStudy (anonymous):

(: okay

OpenStudy (anonymous):

what you think i made it up?

OpenStudy (anonymous):

How many solutions are there for the equation? x2 + 9x = -7 A. 2 B. 1 C. 0 I put A. right? and idk lol this person gave me random answers before just to get medals... dont see the good in that but thats what they did! :P I believe you though (:

OpenStudy (anonymous):

i think you are supposed to write \[x^2+9x+7=0\] and since this doesn't factor over the integers you have to compute \[9^2-4\times 1\times 7\] and see if it is positive, which it is, so you have two answers \

OpenStudy (anonymous):

random answers is not good;(

OpenStudy (anonymous):

Hmmm So then i got it right ... eh?

OpenStudy (anonymous):

yes it is right. two anwers.

OpenStudy (anonymous):

next?

OpenStudy (anonymous):

(: haha woo hoo!! so far so good!!

OpenStudy (anonymous):

A boy standing on a 64 foot cliff drops a coin off the cliff. If the equation of the height h of the coin t seconds after it drops is h = -16t2 + 64, about how many seconds will the coin take to reach the bottom of the cliff? A. 1 second B. 1.5 seconds C. 2 seconds D. 2.5 seconds i guessed on this one and i got B

OpenStudy (anonymous):

\[-16t^2+64=0\] \[16t^2=64\] \[t^2=4\] \[t=2\]

OpenStudy (anonymous):

no guessing.

OpenStudy (anonymous):

^_^ haha okay. Thanks!! (: next one Solve the equation by graphing. x2 + 5x + 8 = 0 A. {-5, 8} B. {2, 4} C. {-4, 5} D. no solution I got D correct? (:

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sits above the x - axis. not zeros

OpenStudy (anonymous):

(: Ivys kinda smart isnt she? lmao haha I guess i remembered more than i thought on this one :P

OpenStudy (anonymous):

So far only one wrong!! ^^

OpenStudy (anonymous):

good. just guessing or doing any work?

OpenStudy (anonymous):

Last one!! Solve the equation by graphing. x2 - 18x + 81 = 0 A. {-2, 9} B. {9/2, 3} C. {9} D. no solution I got C.. am i right? And i have been doing work (:

OpenStudy (anonymous):

\[x^2-18x+81=(x-9)^2\]

OpenStudy (anonymous):

answer is 9

OpenStudy (anonymous):

Woooo!!!!! :D turns out doing the work works mucho bettah

OpenStudy (anonymous):

is there more or is that all?

OpenStudy (anonymous):

yeah go figure nothing like actually computing!

OpenStudy (anonymous):

Nope. I turned it in and i got 100% i could kiss you!! lol I love this cite!!

OpenStudy (anonymous):

done! now forget math and do something else.

OpenStudy (anonymous):

congrats

OpenStudy (anonymous):

Haha i cant :P

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Too much work to do. Too little time.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

ok i will do something else. have fun

OpenStudy (anonymous):

I HAVE to get my grade up to B i HAVE to!!!!!!!!! I about ready to pull out my hair!!! All i know is i AM getting that grade up before i take the exam!!!!!!!! AHHHH!! lol